40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Wed Jan 16 22:07:44 2019, awade, anchal, DailyProgress, PLL, PLL noise from the Marconi 2023A FG IMG_4875.JPGIMG_4877.JPG
    Reply  Fri Jan 18 13:10:02 2019, anchal, DailyProgress, PLL, PLL noise from the Marconi 2023A FG MarconiPLLNoiseAnalysis.pdfMarconi_PLL_Noise.zip
Message ID: 2284     Entry time: Wed Jan 16 22:07:44 2019     Reply to this: 2286
Author: awade, anchal 
Type: DailyProgress 
Category: PLL 
Subject: PLL noise from the Marconi 2023A FG 

[awade, anchal]

This is not a detailed post as I didn't collect data.  We need to check the actual noise performance of the Marconi and compare.  It seems like the marconi used in the PLL has some nasty frequency spikes and generally higher noise. 

Today I dug out a Wenzel crystal OCOX oscillator (Freq 24.483 MHz) from the QIL lab to check the performance of the Marconi FG against a stable reference. These crystals are a good reference to compare to as they should have a noise floor of -165 dBc/Hz vs the marconi performance of -124 dBc/Hz (@ 20 kHz) operating at 470 MHz (datasheet didn't have values for smaller freq).  

The output is +13 dBm so I attenuated by 13 dB to give a simulated BN signal of about 0 dBm (gives mixer gain 0.210V/rad).  This was then substituted for the beat note signal and and the PLL was locked with the Marconi carrier close to the nominal crystal frequency.  Marconi slope was set to 1 kHz/V, LO carrier was +13 dB into mixer (lv13), SR560 gain was 20.  Screen shot picture (attached below) shows the original Marconi (labeled #539) in the reference trace and the spare PSL lab Marconi (labeled BD90200).  Sorry there is no data capture, I was using the SR785 from the EEshop that isn't configured for our lab IP addresses.

What we see is at least 10 dB worse performance for exactly the same marconi settings.  We can also see that there is also a huge spike at about ~1kHz.  This could be the source of some of the humping in the BN spectrum. Not sure about this but maybe some of the LF noise in the spectrum gets convoluted about this strong spike, smearing noise out at the higher frequencies.

The Marconi in the cryo lab, from the cryo-CTN experiment was also procured.  The noise on that Marconi was even lower and did not have any spike artifacts. However, I was able to see a bit of a spike when the external modulation voltage was a little bit negative. 

I then messed around with some of the settings buried in the menu, I turned the external RF clock reference off on the cryo Marconi and also activated the RFDC null with the modulation BNC 50Ω terminated.  After I had changed some of these things I was not quite sure if I recovered the same noise performance from the better units.  Also, I noticed that the 10 kHz/V range on the cryo Marconi didn't want to lock at all, it was as if there was no actuation going on; it worked fine on settings <=4 kHz/V, just not higher, not sure what is wrong here.

We need to do some detailed measurements tomorrow at a range of different frequencies to nail down exactly what is going on here.  As the VCO noise in the PLL loop is not suppressed by the loop gain, this goes strait to the bottom line of the BN spectrum.  We need to know that PLL is not limited by the Marconi noise.

Measurements to do:

  1. PLL BN spectrum with all three Marconis at 24.483 MHz for 0.5 kHz, 1 kHz, 5 kHz and 10 kHz
  2. PLL BN spectrum with all three Marconis at doubled/ tripled and quadrupled frequencies (or whatever Wenzel crystals we can get) to check if its just this frequency range
  3. Debug cryo Marconi and see why it won't lock PLL on the 10 kHz range
  4. Stabilize Marconi to external Rb clock from cryo lab and see if performance is improved
  5. (optional) run 10 MHz line along back of labs to PSL lab from cryo so we have a GPS stabilized reference to work with (low priority)

---

Also, we want to take some BN with the old Marconi swapped out.  You should have all the data you need for drift of AD590 sensor traducers by now. The old TIA circuit is so badly designed that its almost certainly doing a feed forward from room temperature to can heaters. Either install the new circuit proto or turn the vacuum can thermal controls OFF completely. Just install what you have and use the secondary circuit as a your tester.  But don't spend any more time making everything perfect: you are getting 1/t returns on your time on this one. You'll want to box it up and make a yellow Styrofoam insulator box to shield it more from room temp changes.

Attachment 2: IMG_4877.JPG  2.247 MB  | Hide | Hide all
IMG_4877.JPG
ELOG V3.1.3-